OpenStudy (anonymous):

WILL GIVE A MEDAL AND A FAN Find the length of the diameter of a circle whose endpoints are at (11, 4) and (10, 5). A. 1 B. square root 2 C. 2 square root 2

OpenStudy (amorfide):

you have an equation for a circle \[(x+k)^{2}+(y+k)^{2}=r^{2}\] you have the end points of the diameter (11,4) and (10,5) you need to figure out the midpoint of the circle the equation for the midpoint is \[(\frac{ x_{1}+x_{2} }{ 2 } , \frac{ y_{1}+y_{2} }{ 2})\] then we would use the values for x and y, to replace K in the equation of the circle for example if I were to get the midpoint to be (2,1) then my circle equation would be \[(x-2)^{2}+(y-1)^{2}=r^{2}\] now we work out the radius which is the distance of the line equation \[r=\sqrt{(x_{2}-x_{1})^{2} + (y_{2}-y_{1})^{2}}\] then obviously the diameter is double the radius this explanation shows how to write the entire equation of a circle from given points incase you need it in the future